The velocity of sound V in a metal is dependent in the young modulus E and the density e. Use the method of dimension to derive the relationship between the parameters
1
Expert's answer
2021-10-10T16:04:20-0400

Dimensions of young modulus:


"E=[M]^1[L]^{-1}[T]^{-2}"

Dimensions of density:


"\rho=[M]^1[L]^{-3}"

Speed of sound:


"v=[L][T]^{-1}"

Obtain the relationship:

"[L]^1[T]^{-1}=([M]^1[L]^{-1}[T]^{-2})^a·([M]^1[L]^{-3})^b,\\
[L]^1[T]^{-1}=[M]^{a+b}[L]^{-a-3b}[T]^{-2a}."

Equate the power of each dimension (powers on the left with powers on the right):


"[T]^{-1}=[T]^{-2a},\\
-1=-2a→a=\frac12.\\\space\\
[M]^0=[M]^{a+b},\\
0=a+b→b=-\frac12."

Therefore, we have

"\Big([L]^1[T]^{-1}\Big)=\\=\Big([M]^1[L]^{-1}[T]^{-2}\Big)^{\frac12}·\Big([M]^1[L]^{-3}\Big)^{-\frac12},"

which corresponds to


"v=\sqrt\frac E\rho."